# Automorphic vector-forms using the Cohn-Elkies magic functions

### v1

@@ Abstract
In this study, we introduce the theory of what we call Hecke vector-forms. A Hecke vector-form can be viewed as a vector function representation of some quasiautomorphic form that transforms like an automorphic form on an arbitrarily chosen Hecke triangle group. In other words, because quasiautomorphic forms have complicated transformation behavior when compared with automorphic forms, the construction of a Hecke vector-form is to retrieve a transformation behavior analogous to the simpler, automorphic case. In this way, a Hecke vector-form can be viewed as the vector function analogue of an automorphic form. Since our work is for any quasiautomorphic form over an arbitrary Hecke triangle group, we briefly review the construction of such groups. Furthermore, we review the derivation of the hauptmodul, the automorphic forms, and the normalized quasiautomorphic form of weight $2$ for any Hecke triangle group. We then proceed to the theory of Hecke vector-forms and establish the desired transformation behavior with respect to the generators of the associated group. A proof of this fact is strictly elementary, relying on fine properties of binomial coefficients. Lastly, we relate the vector-forms to Hecke automorphic linear differential equations, which are analogues of the frequently researched modular linear differential equations. Our results include Hecke vector-forms of the classical quasimodular forms as the simplest case.
@@ 1 Introduction
Modular forms are meromorphic functions defined on the upper half plane $\mathbb{H}$ with an especially elegant transformation behavior for elements of the modular group
@@
Functions having this transformation behavior on other subgroups of $\mathbf{PSL}_{2}(\mathbb{R})$ are called automorphic forms. Here forward, discussion of Hecke automorphic forms will include the modular forms–after all, the modular group is indeed the “first” Hecke triangle group.
@@
Automorphic forms are an old study; quasiautomorphic forms, however, were defined much more recently (in 1995, see [23]). The span of time between the introduction of automorphic forms and quasiautomorphic forms might give the impression that these latter functions are exotic or only of special interest. It is true, for example, that a quasiautomorphic form does not exhibit the elegant transformation property that is so characteristic of an automorphic form. This makes quasiautomorphic forms more cumbersome in practice, the complication increasing as the depth of the vector space increases. It is perhaps this reason that the functions were ignored for so long. Nonetheless, it was first seen in the breakthrough result [38] (similarly, see [11] and [14]) that the transformation behavior of a quasiautomorphic form is of great power in itself and so here we accept this behavior as basic. Quasiautomorphic forms are the Cohn-Elkies magic functions of mathematical folklore.
@@
The aim of this article is to introduce what we call the Hecke vector-forms (see definition 74), which are special vector valued functions over the Hecke triangle groups. In contrast to other studies like [4] and [15] that combine automorphic forms and vector functions, we aimed here to begin instead with quasiautomorphic forms. In other words, the component functions of the Hecke vector-forms are strictly quasiautomorphic forms. In doing this, we want to provide a cohesive formalism both for working with such functions and also setting the groundwork for a uniform future study. In this spirit, we also prove some results about Hecke vector-forms that have a more or less fundamental feel. These results include a “corrected” transformation formula (see Theorem 76). We also show how vector-forms are deeply connected to a class of automorphic linear differential equations (see Theorem 85).
@@
Our project was not completed in the greatest generality possible (with respect to what is known about quasiautomorphic forms on Schwarz triangle groups), but instead we chose to highlight the very close formal relationship between automorphic forms on the Hecke triangle groups, generally, and the simplest, most familiar instance, the modular forms on the full modular group. Being more precise, we show that with respect to the formalism we introduce, moving between different Hecke triangle groups is simply a matter of accounting for two parameter values: in our notation, these real parameters are $C_{\mathfrak{t}_{\mu}}$ and $\varpi_{\mu}$, a structure constant and a shift, respectively. In this way, our study bears heavily on the classical theory of modular forms. As such, we hope the developments will be useful to the usual practitioners from number theory and physics. The aforementioned practitioners may like to know that there were a few books that were indispensable in making precise many of the ideas that are left out in the more concise research articles. We owe these texts a special mention: [19], [20], [29] , [35], and [39] were never far out of reach.
@@ 2.1 Elementary theory of Hecke triangle groups
  1. 1.
    >> or
  2. 2.
    == or
  3. 3.
@@
Once $\mu$ is chosen, we want a more convenient normalization of the triangle groups for use with quasiautomorphic forms. Toward this end, for a triangle group $(\mu_{1},\mu_{2},\infty)$ let the corner vertices $r_{1},r_{2},$ and $r_{3}$ of the initial curvilinear triangle satisfy
@@
Let $3\leq\mu\in\mathbb{N}$. We denote the normalized $(\mu-2)$-th normalized Hecke triangle group by
@@
Remark 5.
The notation $\mathfrak{H}(\varpi_{\mu})$ is intended to combine the signature $\mathfrak{t}_{\mu}=(2,\mu,\infty)$, the term $\varpi_{\mu}$ from the generator $T_{\mathfrak{t}_{\mu}}z=z+\varpi_{\mathfrak{t}_{\mu}}$, and the normalization found at (5). In particular, while working with Hecke vector-forms, the term $\varpi_{\mu}$ is vital to have on hand.
@@
In Table 1, we give a summary of important features for the first several Hecke triangle groups.
@@
2.2 Darboux-Halphen system gives the Hecke hauptmodul…
@@
We assume from this point forward that $\mathfrak{t}_{\mu}=\mathfrak{H}(\varpi_{\mu})$ (i.e. $\mathfrak{t}_{\mu}$ is normalized). This particular process of finding quasiautomorphic forms amounts to finding solutions of a special differential system that is very much of interest in its own right. These solutions are necessarily quasiautomorphic forms (as shown in Theorem 4 (i) of [9]); however, basic operations on the solutions give us the so-called hauptmodul, the so-called automorphic discriminant, and the automorphic forms analogous to Eisenstein series. We will also find explicitly from the differential solutions the unique weight $2$ normalized quasiautomorphic form. This section comes adapted from [9], essentially with only small changes in notation; this paper is required reading from the perspective of quasiautomorphic forms on triangle groups.
@@
Remark 8.
Let us consider a line of $\mathsf{DH}_{\mathfrak{t}}$ say, for instance
$\theta y_{1}=(a_{\mathfrak{t}}-1)[y_{1}(y_{2}+y_{3})-y_{2}y_{3}]+(b_{\mathfrak{t}}+c_{\mathfrak{t}}-1)y_{1}^{2}.$
Notice that we could consider this in isolation as a differential equation
$\theta y=y^{2}+b_{1}y+b_{0},$
$b_{1}(q)=y_{2}(q)+y_{3}(q)$ and $b_{0}(q)=y_{2}(q)\,y_{3}(q)$ (having ignored the constants for simplicity); this clearly takes the form of a classical Riccati equation. Thus, $\mathsf{DH}_{\mathfrak{t}}$ is an instance of a Riccati system. For the relevance to our study, see equation (1) and the discussion thereof. Also, by [9], the solutions $y$ of system (7) are quasiautomorphic forms of weight $2$.
@@
Remark 14.
Definition 12 is a very deep and interesting result proven as a theorem in [9]. In particular, it says that the hauptmodul of $\mathfrak{H}(\varpi_{\mu})$ is an affine-ratio of differential solutions from $\mathsf{DH}_{\mathfrak{t}_{\mu}}$ (though it holds for more general triangle groups as well, see [9]). It is a generalization of a classical result about the so-called modular $\lambda$-function or the hauptmodul of the subgroup of $\mathfrak{H}(1)$, denoted $\Gamma(2)$ (congruence subgroup of level $2$). Another well-known place $\lambda$ appears is the Jacobi differential equation
$(y^{\prime})^{2}=a^{2}(1-y^{2})(1-\lambda^{2}y^{2})$
(see [19], p. 690).
@@
Example 15.
We provide a graphical example of two (of countably infinite) fundamental domains in $\mathbb{H}$ for the classical modular case or $\mathfrak{H}(1)$.
[…]
Figure 2: Fundamental domains $D_{1}$, and then $D_{2}$ by application of $S_{\mathfrak{t}_{\mu}}z$, for $z\in D_{1}$
@@
Remark 18.
Using Definition 17, we see that the essence of Definition 12 is
$J_{\mathfrak{t}_{\mu}}=[y_{\mathfrak{t}_{\mu},3},y_{\mathfrak{t}_{\mu},1},y_{\mathfrak{t}_{\mu},2}]$
holds. In other words, using the Darboux-Halphen system, the hauptmodul is a careful parameterization of affine space.
Following [24], p. 129-130, let $z_{i}\in\mathbb{R}$. Then the equations
  1. 1.
    $z_{1},z_{2},z_{3},z_{4}]=[z_{3},z_{4},z_{1},z_{2}]$
  2. 2.
    $z_{1},z_{2},z_{3},z_{4}]+[z_{1},z_{3},z_{2},z_{4}]=-\left[z_{1},1,\frac{2z_{1}}{z_{1}+1},0\right]$
  3. 3.
    $z_{1},z_{2},z_{3},z_{4}][z_{1},z_{2},z_{4},z_{5}]=[z_{1},z_{2},z_{3},z_{5}]$.
characterize the cross-ratio. Using the final equation, the so-called multiplicativity equation, which holds when $z_{i}\in\mathbb{C}$ as well, we can remove the cusp of $J_{{\mathfrak{t}_{\mu}}}$ as in
$J^{*}_{\mathfrak{t}_{\mu}}:=[y_{\mathfrak{t}_{\mu},3}(q),y_{\mathfrak{t}_{\mu},1}(q),y_{\mathfrak{t}_{\mu},2}(q)][y_{\mathfrak{t}_{\mu},3}(q),y_{\mathfrak{t}_{\mu},1}(q),\infty,q]$
$=[y_{\mathfrak{t}_{\mu},3}(q),y_{\mathfrak{t}_{\mu},1}(q),y_{\mathfrak{t}_{\mu},2}(q),q].$
In the language of projective geometry, $J^{*}_{\mathfrak{t}_{\mu}}$ is the absolute coordinate of $q$ relative to the reference $J_{\mathfrak{t}_{\mu}}$ (see [10], p. 59, Theorem 2.9.7).
This completes our introduction of the Hecke hauptmodul and the brief review of projective geometry. For a great reference on the latter subject see [10].
2.3 …Darboux-Halphen gives the automorphic forms and the quasiautomorphic forms
@@
Let $3\leq\mu\in\mathbb{Z}$. With some loss of generality, let $w=4k$ be the weight of a function in the space $\mathcal{M}_{w}(\mathfrak{H}(\varpi_{\mu}))$ and $\lfloor x\rfloor$ is the so-called floor function. Then the relation
@@
Quasiautomorphic forms for $\mathfrak{H}(1)$ were introduced in [23] as a vector space generalization of modular forms. A self-contained introduction to these functions can be found in [34].
@@
Remark 32.
We give a short summary in a Figure 3 of the steps behind construction of our desired functions.
[…]
Figure 3: Determination of desired maps on $\mathfrak{H}(\varpi_{\mu})$.
It is also clear that $\Delta^{\prime}_{\mathfrak{t}_{\mu}}/\Delta_{\mathfrak{t}_{\mu}}=E_{\mathfrak{t}_{\mu},2}$ and the Ramanujan differential system gives us the other Eisenstein series; however, we do not indicate these relations in our figure.
@@ 3.2 Required matrix theory
We do not need to laboriously review the following standard matrix operations:
@@
This so-called exchange matrix (as it is given in [21]) is perhaps not typical, but it is important for us. The standard identity $I_{r}$ can be recovered with $\iota_{r}$, using operations we introduce later (as we illustrate in Example 47).
@@
To show the simplicity of our exchange operations, observe
@@
holds and
$\begin{pmatrix}\alpha&\cdots&\beta&\gamma\\ &\ddots&\vdots&\vdots\\ &&\delta&\epsilon\\ &&&\zeta\end{pmatrix}^{\emph{\text{Y}}}=\begin{pmatrix}\gamma&\beta&\cdots&\alpha\\ \vdots&\vdots&\iddots&\\ \epsilon&\delta&&\\ \zeta&&&\end{pmatrix}$
holds.
Example 47.
Note that the usual identity $I_{r}$ for matrix multiplication obeys
$I_{r}^{\emph{Y}}=\begin{pmatrix}1^{0}&&&\\ &1^{1}&&\\ &&\ddots&\\ &&&1^{r}\end{pmatrix}^{\emph{Y}}=\iota_{r}.$
In the present study we shall have more use for the special matrix $\iota_{r}$.
@@
$V_{r}^{\vee}(z)=\begin{pmatrix}1&&\\ z&1&\\ \vdots&\vdots&\ddots&\\ z^{r}&z^{r-1}&\cdots&1\end{pmatrix}$
and
$V_{r}^{\wedge}(z)=\begin{pmatrix}1&z&\cdots&z^{r}\\ &1&\cdots&z^{r-1}\\ &&\ddots&\vdots\\ &&&1\end{pmatrix}$,
etc. With this notation, also observe that $I_{r}$ can be written as either an upper triangular or lower triangular matrix, namely the relation
@@
Example 52.
We point out the following trivial identity that will appear again in the Fundamental Theorem: the relation
$\mathcal{V}_{r}(f_{0},f_{1},\cdots,f_{r})\odot\iota_{r}=\begin{pmatrix}&&&f_{0}^{r}\\ &&f_{1}^{r-1}&\\ &\iddots&&\\ 1&&&\end{pmatrix}$
holds. Therefore, if $r$ is even, then
$\mathcal{V}_{r}(-1,-1,\cdots,-1)=\begin{pmatrix}&&&(-1)^{r}\\ &&(-1)^{r-1}&\\ &\iddots&&\\ 1&&&\end{pmatrix}=\begin{pmatrix}&&&1\\ &&-1&\\ &\iddots&&\\ (-1)^{r}&&&\end{pmatrix}$
holds. If $r$ is odd, then
$-\mathcal{V}_{r}(-1,-1,\cdots,-1)\odot\iota_{r}=\begin{pmatrix}&&&1\\ &&-1&\\ &\iddots&&\\ (-1)^{r}&&&\end{pmatrix}$
holds. These are alternating exchange matrices.
Notation 53.
We introduce shortened notation for the alternating exchange matrices from the previous Example. We write
@@ 3.3 Component functions $g_{U,\ell}$ and $f_{U,k}$
3.4 Hecke vector-forms and transformation formula
@@
Example 73.
Suppose that $\vec{F}_{U}(z)$ and $\vec{G}_{V}(z)$ are two Hecke vector-forms of quasiautomorphic forms $U_{\mathfrak{t}_{\mu},w,r}$ and $V_{\mathfrak{t}_{\mu},w,r}$, respectively. We recall the meaning of notation: $U_{\mathfrak{t}_{\mu},w,r}$ is a quasiautomorphic form on the Hecke triangle group $\mathfrak{t}_{\mu}$ of weight $w$ and depth $r$, etc. Then it follows that
$\vec{F}_{U}(z)=\begin{pmatrix}f_{0}(z)&f_{1}(z)&\cdots&f_{r}(z)\end{pmatrix}^{\text{\emph{T}}}$
and
$\vec{G}_{V}(z)=\begin{pmatrix}g_{0}(z)&g_{1}(z)&\cdots&g_{r}(z)\end{pmatrix}^{\text{\emph{T}}}$
where details of the component functions $f_{i}$ and $g_{k}$ we ignore as they are not essential in this example (and all details can be found in our paper). What is significant, however, is the dot product defined for vectors $\mathbf{x}$ and $\mathbf{y}$ is
$\langle\mathbf{x},\mathbf{y}\rangle:=\mathbf{x}^{\text{\emph{T}}}\mathbf{y}$
where $^{\text{\emph{T}}}$ is the vector transpose operator. Thus, the theory of Hecke vector-forms comes equipped with a simple point-wise inner product.
@@
$\vec{F}_{U}(z,1)$ $=$ $\begin{pmatrix}f_{U,0}(z)&f_{U,1}(z)&\cdots&f_{U,r}(z)\end{pmatrix}^{\mathbf{t}}$
@@
Per usual, $T_{\mathfrak{t}_{\mu}}z$ and $S_{\mathfrak{t}_{\mu}}z$ are generators of $\mathfrak{H}(\varpi_{\mu})$. To recall the matrix definition of $\mathbf{a}_{U}$, see Notation 53. The relations
@@
We prove the statements in order. Firstly, a short calculation from Defintion 74 says
@@
is true, giving us what we wanted per Notation 75.
@@
holds. Taking $\pm\mathbf{a}_{U}$ whether $r$ is even or odd as in Example 52, then we are finished. ∎
Remark 77.
It is instructive to compare Theorem 76, alongside Proposition 57 and Definition 74. In a sense, the Fundamental Theorem can be viewed as variation of the binomial theorem (Proposition 57), but with a parameter $\varpi$ to simplify shifts $z+\varpi$ in the argument of $V^{\vee}_{r}(z)$. Clearly, the vector of Definition 74, namely $G_{U}(z)$, is much more complicated than $B_{r}(z)$, so it is remarkable that we can recover a simple identity at all. It seems it is a consequence of the orthogonality properties of the binomial coefficients (see [33] for an exhaustive treatment of such orthogonality properties).

### v2

@@ Abstract
In this study, we introduce the theory of what we call Hecke vector-forms. A Hecke vector-form can be viewed as a vector function representation of some quasiautomorphic form that transforms like an automorphic form on an arbitrarily chosen Hecke triangle group. In other words, because quasiautomorphic forms have complicated transformation behavior when compared with automorphic forms, the construction of a Hecke vector-form is to retrieve a transformation behavior analogous to the simpler, automorphic case. In this way, a Hecke vector-form can be viewed as the vector function analogue of an automorphic form. Since our work is for any quasiautomorphic form over an arbitrary Hecke triangle group, we briefly review the construction of such groups. Furthermore, we review the derivation of the hauptmodul, the automorphic forms, and the normalized quasiautomorphic form of weight $2$ for any Hecke triangle group. We then proceed to the theory of Hecke vector-forms and establish the desired transformation behavior with respect to the generators of the associated group. A proof of this fact is strictly elementary, relying on fine properties of binomial coefficients. Lastly, we relate the vector-forms to Hecke automorphic linear differential equations, which are analogues of the commonly researched modular linear differential equations. Our results include Hecke vector-forms of the classical quasimodular forms as the simplest case.
@@ 1 Introduction
Modular forms are holomorphic functions defined on the upper half plane $\mathbb{H}$ with an especially elegant transformation behavior for elements of the modular group
@@
Functions having this transformation behavior on other subgroups of $\mathbf{PSL}_{2}(\mathbb{R})$ are called automorphic forms. Here forward, discussion of automorphic forms will include the modular forms–after all, the modular group is indeed the “first” Hecke triangle group.
@@
Automorphic forms are an old study; quasiautomorphic forms, however, were defined much more recently (in 1995, see [23]). The span of time between the introduction of automorphic forms and quasiautomorphic forms might give the impression that these latter functions are only of special interest. It is true, for example, that a quasiautomorphic form does not exhibit the elegant transformation property that is so characteristic of an automorphic form. This makes quasiautomorphic forms more cumbersome in practice, the complication increasing as the depth of the function increases. Nonetheless, it is seen in the breakthrough result [36] (similarly, see [12] and [15]) that the transformation behavior of a quasiautomorphic form is of great power in itself and so here we accept this behavior as basic. That is, quasiautomorphic forms are the Cohn-Elkies “magic” functions of the mathematical folklore around sphere packing problems [11].
@@
The aim of this article is to introduce what we call the Hecke vector-forms (see definition 65), which are special vector valued functions over the Hecke triangle groups. In contrast to studies such as [4] and [16] that introduce automorphic forms as vector functions more abstractly, we aimed here to begin with quasiautomorphic forms modulo the Hecke triangle groups as the vector components. Since, say, the classical modular group is a Hecke triangle group, these vector functions are in a sense already familiar. We have supplied much detail to provide a cohesive formalism both for working with such functions and also setting the groundwork for a uniform future study. In this spirit, we also prove some results about Hecke vector-forms that have a more or less fundamental feel. These results include a “corrected” transformation formula (see Theorem 67). We also show how vector-forms are deeply connected to a class of automorphic linear differential equations (see Theorem 75).
@@
Our project was not completed in the greatest generality possible (with respect to what is known about quasiautomorphic forms on Schwarz triangle groups), but instead we chose to highlight the very close formal relationship between automorphic forms on the Hecke triangle groups, generally, and the simplest, most familiar instance, the modular forms on the full modular group. Being more precise, we show that with respect to the formalism we introduce, moving between different Hecke triangle groups is simply a matter of accounting for two parameter values: in our notation, these real parameters are $C_{\mathfrak{t}_{\mu}}$ and $\varpi_{\mu}$, a structure constant and a shift, respectively. In this way, our study bears heavily on the classical theory of modular forms. As such, we hope the developments will be useful to the usual practitioners from number theory and physics.
@@ 2.1 Elementary theory of Hecke triangle groups
$>\quad\text{or}\quad=\quad\text{or}\quad<,$
@@
Once $\mu$ is chosen, we want a more convenient normalization of the triangle groups for use with quasiautomorphic forms. Toward this end for a triangle group $(\mu_{1},\mu_{2},\infty)$ let the corner vertices $r_{1},r_{2},$ and $r_{3}$ of the initial curvilinear triangle satisfy
@@
Let $3\leq\mu\in\mathbb{N}$. We denote the normalized $(\mu-2)$-th Hecke triangle group by
@@
In Table 1, we give a summary of features for the first several Hecke triangle groups.
@@
2.2 Darboux-Halphen differential systems and the Hecke hauptmodul
@@
We assume from this point forward that $\mathfrak{t}_{\mu}=\mathfrak{H}(\varpi_{\mu})$ (i.e. $\mathfrak{t}_{\mu}$ is normalized). This particular process of finding quasiautomorphic forms amounts to finding solutions of a special differential system that is very much of interest in its own right. These solutions are necessarily quasiautomorphic forms (as shown in Theorem 4 (i) of [9]); however, basic operations on the solutions give us the so-called hauptmodul, the so-called automorphic discriminant, and the automorphic forms analogous to Eisenstein series. We will also find explicitly from the differential solutions the unique weight $2$ normalized quasiautomorphic form. This section comes adapted from [9], basically with only small changes in notation.
@@
This completes our introduction of the Hecke hauptmodul and the brief review of projective geometry. For a complete reference on the latter subject see [10].
2.3 Darboux-Halphen also gives (quasi)automorphic forms
@@
Let $3\leq\mu\in\mathbb{Z}$. Let $w=4k$ be the weight of a function in the space $\mathcal{M}_{w}(\mathfrak{H}(\varpi_{\mu}))$ and $\lfloor x\rfloor$ is the so-called floor function. Then the relation
@@
Quasiautomorphic forms for $\mathfrak{H}(1)$ were introduced in [23] as a generalization of modular forms. A self-contained introduction to these functions can be found in [32].
@@ 3.2 Required matrix theory
The following are all standard matrix operations:
@@
This exchange matrix (as it is called in [21]) is perhaps not typical, but it is important for us.
@@
To illustrate the behavior of our exchange operations, observe
@@
holds, etc.
@@
$V_{r}^{\vee}(z)=\begin{pmatrix}1&&\\ z&1&\\ \vdots&\vdots&\ddots&\\ z^{r}&z^{r-1}&\cdots&1,\end{pmatrix}$
etc. With this notation, also observe that the identity $I_{r}$ can be written as either an upper triangular or lower triangular matrix i.e. the relation
@@
Notation 45.
We introduce shortened notation for the alternating exchange matrix. We write
@@ 3.3 Component functions $g_{U,\ell}$ and $f_{U,k}$
3.4 Transformation formulas of a Hecke vector-form
@@
$\vec{F}_{U}(z,1)$ $=$ $\begin{pmatrix}f_{U,0}(z)&f_{U,1}(z)&\cdots&f_{U,r}(z)\end{pmatrix}^{\text{\emph{T}}}$
@@
Per usual, $T_{\mathfrak{t}_{\mu}}z$ and $S_{\mathfrak{t}_{\mu}}z$ are generators of $\mathfrak{H}(\varpi_{\mu})$. To recall the matrix definition of $\mathbf{a}_{U}$, see Notation 45. The relations
@@
We prove the statements in order. Firstly, a short calculation from Defintion 65 says
@@
is true, giving us what we wanted per Notation 66.
@@
holds. Taking $\pm\mathbf{a}_{U}$ whether $r$ is even or odd as in Example LABEL:exa:altExchange, then we are finished. ∎